# RLHFSpec: Breaking the Efficiency Bottleneck in RLHF Training via Adaptive Drafting

### v1

@@ Abstract.
Reinforcement Learning from Human Feedback (RLHF) is an important fine-tuning technique for large language models (LLMs) and comprises three stages: generation, inference, and training. The generation stage generates samples that are then used to infer learnable experiences for training. We observe that the generation stage is the bottleneck of the entire execution process and consider it a key point for optimization. Specifically, we realize the first attempt to integrate speculative decoding into the RLHF generation stage and propose RLHFSpec, an RLHF system that accelerates generation execution with adaptive speculative decoding and sample reallocation. To fully exploit the performance potential provided by speculative decoding, especially dealing with the dynamic workload of the generation stage, RLHFSpec proposes a workload-aware drafting strategy selection mechanism, which selects the near-optimal strategy by jointly considering the verification cost and the number of accepted tokens. Moreover, RLHFSpec also proposes sample reallocation to fully utilize the GPU resources, and optimizes it with an efficient sample migration mechanism. The experimental results show that the RLHFSpec can achieve higher throughput in the generation stage compared to state-of-the-art works. Moreover, due to the effective alleviation of the generation bottleneck, RLHFSpec also shows significant performance speedup in the entire RLHF execution.
@@ 1. Introduction
  • We develop an RLHF system RLHFSpec that accelerates RLHF generation with adaptive speculative decoding and sample reallocation. We perform a thorough experimental comparison with existing systems and demonstrate significant performance improvements over the state-of-the-art works.
@@ 4. Design Overview
To leverage the observations discussed above, we propose RLHFSpec, an RLHF system that accelerates RLHF generation with adaptive speculative decoding and sample reallocation. The RLHFSpec system is composed of two major components, including workload-aware drafting strategy selector (Section 5) and lightweight sample reallocator (Section 6). The workflow of the RLHFSpec system is shown in Figure 6.
@@ 7.7. Overhead Analysis
The overhead introduced by RLHFSpec comes from three parts, including workload-aware drafting strategy selection ($WDS$), sample reallocation decision ($SRD$), and sample migration ($SM$). WDSWDS predicts the number of accepted tokens and execution time of the tokens in the speculative tree, and traverses the tree to select the near-optimal strategy. $SRD$ determines the reallocation policy based on the workload of each instance. $SM$ migrates the KVCache of samples across instances. The overhead incurred by $WDS$, SRDSRD, and $SM$ is less than 1.74% of the total execution time, which imposes little impact on generation execution. There are also offline overhead associated with data collection to train the predictors and determine the throughput turning points. This overhead totals approximately 15 minutes and is a one-time cost, which is acceptable in comparison to the longer duration of RLHF training.
@@ 8.1. RLHF Training Optimizations
Researchers have proposed various optimizations to accelerate RLHF training (sheng2025hybridflow; hu2024openrlhf; mei2024realhf; zhong2025streamrl; yao2023deepspeed; havrilla2023trlx; xiao2023adaptive). Verl (sheng2025hybridflow) proposes a hierarchical hybrid programming model to enable flexible dataflow representation and efficient execution. OpenRLHF (hu2024openrlhf) places each model on separate devices and designs customized parallel strategies for different tasks. ReaLHF (mei2024realhf) formulates the RLHF execution as an augmented dataflow and uses a search algorithm to discover efficient execution plans. However, these works fundamentally employ autoregressive decoding during the generation stage, leading to inefficient execution and low resource utilization. In contrast, RLHFSpec utilizes the performance potential derived from the application of speculative decoding in generation execution and proposes an adaptive drafting strategy selection mechanism, along with sample reallocation, to fully exploit these opportunities. Through these methods, RLHFSpec overcomes the constraints of autoregressive decoding, which is orthogonal to existing RLHF works.
@@ 9. Conclusion
In this paper, we introduce RLHFSpec, an RLHF system featuring adaptive speculative decoding and sample reallocation. RLHFSpec designs a workload-aware drafting strategy that dynamically selects the drafting strategy for varying generation workloads. RLHFSpec also proposes sample reallocation as the key measure to fully utilize the GPU resources and thus improve the overall performance. To reduce the sample reallocation cost, an efficient two-stage sample migration is employed. Evaluation results show that RLHFSpec delivers higher throughput in RLHF execution.

### v2

@@ Abstract.
Reinforcement Learning from Human Feedback (RLHF) is an important fine-tuning technique for large language models (LLMs) and comprises three stages: generation, inference, and training. The generation stage generates samples that are then used to infer learnable experiences for training. We observe that the generation stage is the bottleneck of the entire execution process and consider it a key point for optimization. Specifically, we realize the first attempt to integrate speculative decoding into the RLHF generation stage and propose RLHFSpec, an RLHF system that accelerates generation execution with efficient speculative decoding and sample reallocation. To fully exploit the performance potential provided by speculative decoding, especially dealing with the dynamic workload of the generation stage, RLHFSpec proposes a workload-aware drafting strategy selection mechanism, which selects the near-optimal strategy by jointly considering the verification cost and the number of accepted tokens. Moreover, RLHFSpec also proposes sample reallocation to fully utilize the GPU resources, and optimizes it with an efficient sample migration mechanism. The experimental results show that the RLHFSpec can achieve higher throughput in the generation stage compared to state-of-the-art works. Moreover, due to the effective alleviation of the generation bottleneck, RLHFSpec also shows significant performance speedup in the entire RLHF execution.
@@ 1. Introduction
  • We develop an RLHF system RLHFSpec that accelerates RLHF generation with efficient speculative decoding and sample reallocation. We perform a thorough experimental comparison with existing systems and demonstrate significant performance improvements over the state-of-the-art works.
@@ 4. Design Overview
To leverage the observations discussed above, we propose RLHFSpec, an RLHF system that accelerates RLHF generation with efficient speculative decoding and sample reallocation. The RLHFSpec system is composed of two major components, including workload-aware drafting strategy selector (Section 5) and lightweight sample reallocator (Section 6). The workflow of the RLHFSpec system is shown in Figure 6.
@@ 7.7. Overhead Analysis
The overhead introduced by RLHFSpec comes from three parts, including workload-aware drafting strategy selection ($WDS$), sample reallocation decision ($SRD$), and sample migration ($SM$). WDSWDS predicts the number of accepted tokens and execution time of the tokens in the speculative tree, and traverses the tree to select the near-optimal strategy. $SRD$ determines the reallocation policy based on the workload of each instance. $SM$ migrates the KVCache of samples across instances. The overhead incurred by $WDS$, SRDSRD, and $SM$ is less than 3.87% of the total execution time, which imposes little impact on generation execution. There are also offline overhead associated with data collection to train the predictors and determine the throughput turning points. This overhead totals approximately 15 minutes and is a one-time cost, which is acceptable in comparison to the longer duration of RLHF training.
@@ 8.1. RLHF Training Optimizations
Researchers have proposed various optimizations to accelerate RLHF training (sheng2025hybridflow; hu2024openrlhf; mei2024realhf; zhong2025streamrl; yao2023deepspeed; havrilla2023trlx; xiao2023adaptive). Verl (sheng2025hybridflow) proposes a hierarchical hybrid programming model to enable flexible dataflow representation and efficient execution. OpenRLHF (hu2024openrlhf) places each model on separate devices and designs customized parallel strategies for different tasks. ReaLHF (mei2024realhf) formulates the RLHF execution as an augmented dataflow and uses a search algorithm to discover efficient execution plans. However, these works fundamentally employ autoregressive decoding during the generation stage, leading to inefficient execution and low resource utilization. In contrast, RLHFSpec utilizes the performance potential derived from the application of speculative decoding in generation execution and proposes an workload-aware drafting strategy selection mechanism, along with sample reallocation, to fully exploit these opportunities. Through these methods, RLHFSpec overcomes the constraints of autoregressive decoding, which is orthogonal to existing RLHF works.
@@ 9. Conclusion
In this paper, we introduce RLHFSpec, an RLHF system featuring efficient speculative decoding and sample reallocation. RLHFSpec designs a workload-aware drafting strategy that dynamically selects the drafting strategy for varying generation workloads. RLHFSpec also proposes sample reallocation as the key measure to fully utilize the GPU resources and thus improve the overall performance. To reduce the sample reallocation cost, an efficient two-stage sample migration is employed. Evaluation results show that RLHFSpec delivers higher throughput in RLHF execution.